OpenStudy (anonymous):

If \[n \sqrt{a}=r\] , then which of the following are true statements? Check all that apply. A. r^n=a B. a^1/n=r C. a^r=n D. n^1/r=a

OpenStudy (anonymous):

A??? o.O

OpenStudy (anonymous):

Well you have: \[\eqalign{ &\sqrt[n]{a}=r \\ &a^{1/n}=r \\ &(a^{1/n})^n=r^n \\ &a^{(n/n)}=r^n \\ &a=r^n }\]

OpenStudy (anonymous):

so only A and B would be the answers?

OpenStudy (anonymous):

@KeithAfasCalcLover

OpenStudy (anonymous):

YEP!

OpenStudy (anonymous):

thank you!! :)

OpenStudy (anonymous):

Anytime ali :)